OpenStudy (brucebaner):

A computer downloads files at a constant rate. The table shows how many megabytes the computer downloads over specific lengths of time. What is the slope of the line that represents the computer downloads? Computer Downloads Minutes 3 5 7 9 Megabytes 11.25 18.75 26.25 33.75

OpenStudy (brucebaner):

@SolomonZelman

OpenStudy (brucebaner):

@SuperflyTyGuy

OpenStudy (brucebaner):

@confluxepic

OpenStudy (brucebaner):

11.25/3=3.75 18.75/5=3.75 26.25/7=3.75 33.75/9=3.75

OpenStudy (brucebaner):

So your slope is 3.75

OpenStudy (brucebaner):

me i right @confluxepic

OpenStudy (confluxepic):

@TheSmartOne @~

OpenStudy (brucebaner):

im i right @SuperflyTyGuy

OpenStudy (superflytyguy):

It's 3.75

OpenStudy (brucebaner):

i know that is why 3.75

OpenStudy (brucebaner):

what im asking is that if im right or not

OpenStudy (superflytyguy):

Yes you're right

OpenStudy (brucebaner):

thx